OpenStudy (anonymous):

There are 10 boys trying out for the track team. Truman and Ryan are the two fastest runners. If the coach randomly picks two boys to run a sprint, what is the probability that he will pick Truman and then Ryan? 1/100 1/90 1/5

OpenStudy (anonymous):

P ( Truman) = 1/10 P ( Ryan) = 1/9 P ( Truman and Ryan) = 1/10 * 1/9 = ....

OpenStudy (anonymous):

1/90